# Leveraging Sub-Optimal Data for Human-in-the-Loop Reinforcement Learning

### v1

@@ 5 Experiments
This section considers the following five research questions:
  • In the low feedback regime, can SDP improve upon existing scalar- and preference-based HitL RL methods?
  • Can SDP effectively leverage sub-optimal data from tasks other than the target task?
  • What is the contribution of each phase of SDP?
  • How does the amount of sub-optimal data affect the performance of SDP?
  • How does the amount of feedback affect the performance of SDP?
@@ 5.1 Experimental Design
To demonstrate the generality and effectiveness of SDP, we apply SDP to (1) scalar-based and (2) preference-based HitL RL approaches. For the scalar-based experiments, we combine SDP with R-PEBBLE (a regression variant of PEBBLE Lee et al. (2021)). We compare SDP + R-PEBBLE against the following benchmarks: R-PEBBLE, Deep TAMER Warnell et al. (2018) (a scalar feedback HitL RL algorithm), and SAC. For the preference-based experiments, we further show the robustness of SDP by combining it with three SOTA preference-based algorithms. This results in the following benchmarks: PEBBLE Lee et al. (2021), RUNE Liang et al. (2022), SURF Park et al. (2022), and SAC. We treat SAC Haarnoja et al. (2018) as an oracle baseline because it learns while accessing the true reward function, which is unavailable to the other algorithms (SAC is the core RL algorithm used across all baselines). We include it as a baseline only to show an upper bound of task performance. See Appendix A for full implementation and hyperparameter details.
For evaluation, we show average offline performance (i.e., freeze the policy and evaluate it with no exploration) over five episodes using either the ground truth reward function (DMControl experiments) or the success rate (Meta-world experiments). We perform this evaluation every 10,000 training steps. To systemically evaluate performance, we use a scripted teacher that provides either a scalar rating of a single trajectory segment or preferences between two trajectory segments according to the true reward function. To thoroughly test the effectiveness of SDP, we perform evaluations on three robotic locomotion tasks from the DMControl Suite: Walker-walk, Cheetah-run, and Quadruped-walk, and six robotic manipulation tasks from Meta-world: Hammer, Door-unlock, Door-lock, Drawer-open, Door-open, and Window-open.
In our experiments, the results are averaged over five seeds with any shaded regions or error bars indicating 95% confidence intervals. To test for significant differences in final performance and learning efficiency (e.g., area under the curve, AUC), we perform Welch t-tests (equal variances not assumed) with a p-value of 0.05. See Appendix B, Tables 4, and 5 for a summary of final performance and AUC across all tasks.
@@ 5.2 Locomotion and Manipulation Results
We first evaluate the utility of SDP in the scalar-based HitL RL setting. Across all tested DMControl environments, we found that SDP (purple curve) significantly improves either the final performance or the learning efficiency (e.g., AUC) compared to R-PEBBLE (navy curve) and Deep TAMER (yellow curve) (see Figure 2). More impressively, we found that SDP achieves comparable performance to SAC (red curve), which uses the ground truth reward function, using as little as 60 feedback queries.
Preference Feedback Experiments
To further showcase the effectiveness of SDP, we perform preference learning experiments in both DMControl and Metaworld suites. Considering all three preference-based algorithms in the nine environments, SDP significantly (p $<$ 0.05) improved learning (i.e., either final performance or AUC) in 15 out of the 27 experiments (see Figure 3 and Table 4 in Appendix B). In the remaining experiments, there were no significant differences in the performance between SDP and the baseline algorithms. The addition of SDP (i.e., SDP + base algorithm) never hurt performance. Furthermore, we performed additional experiments using an imperfect teacher and the results are consistent (see Figure 11 in Appendix B).
Experiments with Sub-Optimal data from Different Tasks
Our earlier experiments demonstrated that if we have sub-optimal data for our task of interest (i.e., target task), then SDP can leverage this data to improve reward learning HitL RL methods. However, in many cases, we may also have sub-optimal data from other related tasks. This section considers whether SDP can improve performance on a target task if it leverages sub-optimal data from a different task.
@@
We perform three preference learning experiments, comparing PEBBLE with SDP + PEBBLE, using sub-optimal data from a different prior task that has the same virtual robot (i.e., only reward function differs): (1) Walker-stand for the target task of Walker-walk, (2) Quadruped-walk for the target task of Quadruped-run, and (3) Drawer-open for the target task of Door-open. To obtain the sub-optimal data for the prior tasks, we gathered transitions from partially trained policies as opposed to using random policies. This ensured that the distribution of sub-optimal data differed between the prior and target tasks. See Appendix A.2 for further details on the experiment setup. Figure 4a demonstrates the flexibility of SDP as it can successfully leverage sub-optimal data from related tasks (green curve) and still outperform PEBBLE (blue curve) in two out of three of the environments.
(a) These figures highlight that SDP can leverage sub-optimal data from tasks other than the target task and still outperform PEBBLE.
(b) These figures demonstrate that the complete SDP outlined in Algorithm 1 achieves significantly greater performance than various SDP variants.
(c) These figures highlight that SDP is effective with differing feedback budgets but is more successful with larger amounts of prior data.
Figure 4: This figure showcases several SDP ablations: SDP transfer ablations (Top), SDP component ablations (Middle), and amount of prior pre-training data and feedback ablations (Bottom).
5.3 Ablation Studies
To further understand the effectiveness of SDP, we perform further analysis of SDP across three dimensions: (1) the phases of SDP, (2) the amount of sub-optimal data, and (3) the number of feedback queries. To cover scalar- and preference-based HitL RL settings in both DMControl and Meta-world, we consider SDP + R-PEBBLE (i.e., scalar feedback) in Walker-walk and SDP + PEBBLE (i.e., preference feedback) in Cheetah-run and Door-open.
@@ 5.3 Ablation Studies
To understand the effect of each component of SDP, we perform three ablations. First, we evaluate the effect of each phase of SDP individually (i.e., the reward model pre-train phase and the agent update phase). Figure 4b–leftmost demonstrates the importance of using both phases in SDP for scalar-based HitL RL approaches. We found that the SDP variants that only use one of the phases (green and gray curves) result in worse performance than the full SDP (purple curve). We observed similar results in the Cheetah-run and Door-open preference learning experiments (see Figures 8a and 8b in Appendix B).
Second, in the reward model pre-training phase, the goal is for the reward model to learn to output zero. However, a trivial means to achieve an output of zero is to set all weights and biases in the neural network to zero. Therefore, we compare the full SDP to SDP using a zero-weight initialization as a replacement for the reward model pre-training. We found that only using a zero-weight initialization for the reward model instead of the pre-training phase results in significantly degraded performance (see the green curve in Figure 4b–rightmost). This is not surprising, as previous works have found that a zero-weight initialization can negatively affect the training of neural networks Blumenfeld et al. (2020); Zhao et al. (2022). Figure 7 in Appendix B further shows that the reward model pre-training phase does not result in zero-weight values for the reward model. This explains why our pre-training does not lead to the same poor performance as the zero-weight initialization.
[…]
@@
[…]
Effect of Feedback Amount
To investigate the effect of the amount of feedback on the performance of SDP, we evaluated SDP and PEBBLE (and their scalar feedback variants) using three different feedback budgets. For both Cheetah-run and Walker-walk environments, we considered budgets $N\in\{60,100,200\}$. In Figure 4c (middle and right-most), we show that SDP (purple curves) consistently outperforms PEBBLE (blue curves) and R-PEBBLE (navy curves). This further supports SDP’s effectiveness across a range of feedback queries. We achieved similar results in the Door-open environment (see Figure 9c in Appendix B).
@@ 6 Conclusion
In this work, we present SDP, an approach that improves the feedback efficiency for HitL RL algorithms. SDP is specifically designed to leverage reward-free, sub-optimal data for scalar- and preference-based HitL RL approaches. By pseudo-labeling low-quality data with zero, we secure “free” reward labels to pre-train the reward model, giving the reward model a head start in learning. Through this process, the pre-trained reward model can identify that low-quality transitions should have a low reward value, all without having acquired any actual feedback. Our extensive experiments in both DMControl and Metaworld suites demonstrate that SDP can significantly improve both preference- and scalar-based reward learning algorithms. This work takes an important step towards considering how sub-optimal data can be leveraged for HitL RL. Future work should consider other mechanisms for leveraging low-quality data. One interesting direction could be using sub-optimal data from multiple tasks to improve HitL RL algorithms. Another possibility is to consider approaches that can leverage both sub-optimal and expert data.
@@
Appendix A Experimental Details
@@ A.2 Training details
Furthermore, we use most of the existing reward model hyperparameters used in PEBBLE, however, we adjusted the following four hyperparameters: feedback frequency, amount of feedback per session, trajectory segment size (only for Meta-world), and activation function for the final NN layer. We adjusted the first two hyperparameters because PEBBLE originally used a significantly larger feedback budget, therefore we wanted the feedback schedule to better reflect a smaller feedback budget. We used a different trajectory segment size for Meta-world because we wanted to keep the segment sizes the same across both the DMControl and Meta-world environments. Moreover, we found that the output activation function could significantly affect learning, therefore we tested all benchmarks using both Tanh (original activation used) and Leaky-ReLU and chose the reward model that achieved better final performance. For the RUNE and SURF baselines, we use any hyperparameters associated with their specific algorithm according to the original paper (see Table 2). For a fair comparison with SDP, we provide all HiTL baselines (e.g., PEBBLE, R-PEBBLE, Deep TAMER, RUNE, and SURF) with the sub-optimal data set to be used in both the reward model and by the RL agent.
Furthermore, to select trajectory segments for the teacher to provide feedback on, we use uniform sampling in the DMControl tasks and disagreement sampling in the Meta-world tasks. Disagreement sampling is a popular active learning approach in which trajectories with higher uncertainty (based on an ensemble of neural networks) are more likely to be sampled Christiano et al. [2017]. As for the SAC hyperparameters, we use the values found in Table 1.
@@
Initial temperature .1
Discount .99
Batch size
1024 (DMControl)
512 (Metaworld)
Critic, Alpha, Actor:
$\beta_{1},\beta_{2}$
(.9,.999).9.999(.9,.999)( .9 , .999 )
Critic $\tau$ 0.005
@@
Learnable temperature True
Actor and Critic: # of hidden layers
2 (DMControl)
3 (Meta-world)
Actor and Critic: # of hidden units per layer
1024 (DM Control)
256 (Meta-world)
Actor and Critic learning rate
0.00005 (Cheetah-run pref fb exp)
0.0001 (Cheetah-run scalar fb exp)
0.0005 (Walker-walk)
0.0001 (Quadruped)
0.0003 (Meta-world)
Alpha learning rate 0.0001
Actor update frequency 1
Number of training steps
1 million for all experiments except
Walker-walk (Pref and scalar feedback exps.), Door-lock
Window-open
Table 1: Hyperparameters for SAC agent used in all experiments (both preference and scalar feedback variants).
@@
SURF: Threshold $\mu$99
SURF: Lambda $\mu$ 1
SURF: $\mu$ 4
SURF: Inverse label ratio 10
SURF: Data augmentation window 5
SURF: Crop Range 5
RUNE: Beta schedule Linear decay
RUNE: Beta init 0.05
RUNE: Beta decay 0.00001
Table 2: Hyperparameters for SURF and RUNE. All are taken from the original authors’ implementations.
@@
Segment size 50
# of random pre-training steps (i.e., sub-optimal data transitions) 50000
# of unsupervised exploration steps 9000
Include unsup steps as sub-opt data for SDP
All experiments except
Walker-walk (Pref and scalar feedback exps.), Door-open and
Quadruped-walk (Pref learning exps.)
Learning rate 0.0003

### v2

@@ 5 Experiments
This section considers the following four research questions (RQ’s):
  • RQ 1:
    Can SDP improve upon existing scalar- and preference-based RL methods?
  • RQ 2:
    Can SDP effectively leverage sub-optimal trajectories from different tasks to improve performance on a target task?
  • RQ 3:
    Can SDP be used with real human feedback?
  • RQ 4:
    How sensitive is SDP to various hyperparameters?
@@ 5.1 Experimental Design
To demonstrate the versatility of SDP, we apply SDP to both preference and scalar-based RL approaches. However, as preference feedback can be less time-consuming than scalar feedback, we primarily concentrate on preference-based RL in our experiments, exploring scalar feedback in a smaller capacity. For the preference-based experiments, we combine SDP with four contemporary preference-based algorithms: PEBBLE (Lee et al., 2021a), RUNE (Liang et al., 2022), SURF (Park et al., 2022), and MRN (Liu et al., 2022). We benchmark the performance of the four algorithms augmented with SDP against their original versions without SDP, as well as against SAC. We treat SAC (Haarnoja et al., 2018) as an oracle (i.e., upper bound) because it learns while accessing the ground truth reward function, which is unavailable to the other algorithms. For the scalar-based experiments, we combine SDP with R-PEBBLE (a regression variant of PEBBLE). We compare SDP + R-PEBBLE against R-PEBBLE, Deep TAMER (Warnell et al., 2018) (a scalar feedback RL algorithm), and SAC. We note that SAC is the core RL algorithm used across all baselines.
Implementation Details
For SDP, we collected sub-optimal trajectories via a random policy. In particular, we used $50000$ state, action transitions for all experiments in Section 5.2. Note that we do not require explicit access to a sub-optimal policy; we only require state, action transitions from said policy. Moreover, to ensure a fair comparison across algorithms, we maintained equal feedback budgets for all algorithms within each environment, while adjusting the budget across environments to reflect their degree of difficulty. See Appendix A for a complete overview of the implementation process and specific hyperparameters for all algorithms.
Evaluation
We show average offline performance (i.e., freeze the policy and evaluate it with no exploration) over ten episodes using either the ground truth reward function (DMControl experiments) or the success rate (Meta-World experiments). It is important to note that only SAC has access to the ground truth reward function. We perform this evaluation every $10000$ training steps. To systemically evaluate performance, we use a simulated teacher that provides either a scalar rating of a single trajectory segment or preferences between two trajectory segments according to the ground truth reward function. To thoroughly test the effectiveness of SDP, we perform evaluations on four robotic locomotion tasks from the DMControl Suite: Walker-walk, Cheetah-run, Quadruped-walk, and Cartpole-swingup, and five robotic manipulation tasks from Meta-World: Hammer, Door-unlock, Door-lock, Drawer-open, and Window-open. In our experiments, the results are averaged over five seeds with shaded regions or error bars indicating $95$% confidence intervals. To test for significant differences in final performance (i.e., the undiscounted return) and learning efficiency (i.e., the total area under the return curve, AUC), we perform Welch t-tests (equal variances not assumed) with a p-value of $0.05$. See Appendices D.8 and D.9, Tables 9-14 for a summary of final performance and AUC across all experiments.
@@ 5.2 Locomotion and Manipulation Results
Preference Feedback Experiments
We first address RQ 1 by evaluating the utility of SDP in the preference-based RL setting. Considering all four preference-based algorithms in the nine environments, SDP significantly (p $<0.05$) improved learning (i.e., either final performance or AUC) in $23$ out of the $36$ experiments (see Figure 2). In the remaining experiments, there were no significant differences in the performance between SDP and the baseline algorithms. The addition of SDP (i.e., SDP + base algorithm) never statistically hurt performance.
Figure 2: Results from the preference feedback experiments in the DMControl and Meta-World suites show mean AUC $\pm$ 95% confidence intervals. * indicates that SDP + the base preference learning algorithm achieves a statistically greater score than the base preference learning algorithm.
Figure 3: In the scalar feedback experiments on the DMControl environments, SDP significantly outperforms R-PEBBLE and Deep TAMER and achieves comparable performance to SAC.
@@
Continuing our investigation into RQ 1, we now evaluate the performance of SDP in the scalar-based RL setting. We performed evaluations in Walker-walk, Cheetah-run, and Quadruped-walk. In Figure 3, we found that SDP (purple curve) significantly improves either the final performance or AUC compared to R-PEBBLE (navy curve) and Deep TAMER (yellow curve). More impressively, we found that SDP achieves comparable final performance to SAC (red curve), which uses the ground truth reward function, using as little as $60$ feedback queries.
Leveraging Different Task Data in SDP
@@
Our previous experiments showed that SDP can leverage sub-optimal data from the target task to improve human-in-the-loop RL methods. Now, addressing RQ 2, we investigate whether SDP can similarly use sub-optimal data from related tasks to improve performance on the target task. We perform three preference learning experiments, comparing PEBBLE with SDP + PEBBLE, using sub-optimal data from a different prior task that has the same virtual robot: (1) Walker-stand for Walker-walk, (2) Quadruped-walk for Quadruped-run, and (3) Drawer-open for Door-open. To obtain the sub-optimal data for the prior tasks, we gathered transitions from partially trained policies as opposed to using random policies. Each partially trained policy achieved a final score of approximately $15$-$20$% of that achieved by a fully trained policy. This ensured that the distribution of sub-optimal data differed between the prior and target tasks. See Appendix A.2 for further details on the experiment setup. Figure 4 demonstrates that in all three tested environments, SDP can successfully leverage sub-optimal data from related tasks (green curve) as it achieved similar performance to SDP when leveraging target task data (purple curve).
Figure 4: This highlights that SDP can leverage sub-optimal data from different prior tasks as it performed comparable to SDP when using target task sub-optimal data.
5.3 Preference Learning with Human Feedback
To evaluate human-in-the-loop RL algorithms, we argue that it is critical to understand their efficacy with human teachers. However, human user studies do not appear to be widely adopted in the current literature. To empirically investigate this, we conducted a survey of $45$ preference learning studies from 2012 to 2024111See Appendix B.2 for more details. to understand the prevalence of human user studies in the preference learning literature. We found that fewer than $50$% tested their proposed algorithms with human participants who were not also the authors. Moreover, of those studies that did involve human participants, only $41$% included non-expert individuals, while none provided sufficient demographic information about their participants (e.g., gender, race/ethnicity, level of education). These limitations raise significant concerns about whether and how current preference learning algorithms generalize to different populations.
To that end, we now address RQ 3 and evaluate the efficacy of SDP with real human feedback. We conduct an ethics-approved human-subject study of $16$ participants ($9$ male, $7$ female). Age ranges were collected: $18$-$24$ $(6)$, $25$-$30$ $(7)$, and $50$-$70$ $(3)$. Participants self-identified their membership in racial groups: South Asian $(6)$, East Asian $(3)$, White $(4)$, Middle Eastern or North African $(2)$, and multi-racial $(1)$. The participants’ highest educational attainments were: high school diploma $(3)$, Bachelor’s degrees $(8)$, and Master’s degrees $(5)$, and their expertise varied across AI/ML $(12)$, other computer science topics $(1)$, and non-computer science fields $(3)$. This highlights the diversity of our participants in terms of demographics and expertise.
This user study compares the performance of SDP and PEBBLE in two DMControl environments: Pendulum-swingup (with $7$ participants) and Cartpole-swingup (with $12$ participants). We focused our comparison to PEBBLE as it performed comparably to the other preference learning baselines in Section 5.2. We use a between-subjects experimental design — each participant provides preferences for a single seed of both SDP and PEBBLE algorithms. The preference budgets for Pendulum-swingup and Cartpole-swingup were 40 and 48, respectively. We selected these environments specifically because they could be solved with fewer preferences, aiming to reduce the overall time commitment required from participants. Each trial for a single environment lasted approximately 11111.51.51.51.5 hours. See Appendix B for more details including user instructions and interface.
We visualize both algorithms’ final performance and AUC in Figure 5. We found that in both environments SDP (purple plots) maintains significant (p $<0.05$) performance gains over PEBBLE (blue plots) in terms of either final performance or AUC. In Cartpole-swingup, we also observed consistent performance from SDP regardless of whether the teacher was human or simulated. This suggests that our prior results with simulated teachers can generalize to settings where human feedback is provided. Moreover, we observed no significant differences in SDP’s effectiveness across demographic factors, including gender, age, educational, and computer science background (see Appendix C, Tables 7 and 8). This finding is particularly encouraging for the potential real-world deployment of SDP, highlighting its usability for non-expert users across diverse demographics.
Figure 5: This demonstrates that SDP can significantly outperform PEBBLE in terms of both AUC (left) and final performance (right) even when human teachers are providing preferences. * denotes a statistically significant difference between SDP and PEBBLE.
5.4 Ablation and Sensitivity Studies
To further understand the effectiveness of SDP, we perform further analysis of SDP across three dimensions: (1) the phases of SDP, (2) the number of feedback queries, and (3) the amount of sub-optimal data. This analysis aims to address RQ 4 and provide a deeper understanding of the factors influencing SDP’s performance. For these experiments, we focus on SDP + R-PEBBLE in the Walker-walk environment. Additional results for Cheetah-run can be found in Appendix D.5.
@@ 5.4 Ablation and Sensitivity Studies
First, we evaluate the effect of each phase of SDP individually, the reward model pre-train phase, and the agent update phase. Figure 6 (leftmost) demonstrates the importance of using both phases in SDP for scalar-based RL approaches. We found that the SDP variants that only use one of the phases (green and gray curves) result in worse performance than the full SDP (purple curve).
Effect of Feedback Amount
We evaluate SDP and R-PEBBLE with feedback budgets $\in[60,100,200]$ to analyze the impact of feedback quantity on performance. As shown in Figure 6 (middle), SDP (purple curves) consistently outperforms R-PEBBLE (navy curves), further demonstrating its effectiveness across varying feedback levels.
@@
We evaluate the performance of SDP using varying amounts of sub-optimal transitions $\in[5000,15000,50000]$. Figure 6 (rightmost) indicates that while $5000$ transitions (gray curve) led to the poorest performance, increasing this amount to $15000$ or $50000$ (green and purple curves) yielded comparable or improved results, suggesting that more sub-optimal data can benefit SDP.
Figure 6: These figures show ablation and sensitivity studies of SDP in Walker-walk.
@@ 6 Conclusion
In this work, we present SDP, an approach that improves the feedback efficiency for human-in-the-loop RL algorithms. SDP is specifically designed to leverage reward-free, sub-optimal data for scalar- and preference-based RL approaches. By pseudo-labeling low-quality data with the minimum environment rewards, we can pre-train the reward model without the need for human labeling. This provides the reward model with a head start in learning. This head start allows the reward model to learn to associate low-quality transitions with low reward values, even before receiving any actual human feedback. Our simulated teacher experiments in DMControl and Meta-World suites demonstrate that SDP can significantly improve a variety of preference- and scalar-based reward learning algorithms. Importantly, we further validate the real-world applicability of SDP by demonstrating its success in a $16$-person user study. This work takes an important step towards considering how sub-optimal data can be leveraged for human-in-the-loop RL.
Acknowledgments
Part of this work has taken place in the Intelligent Robot Learning Lab at the University of Alberta, which is supported in part by research grants from Alberta Innovates; Alberta Machine Intelligence Institute (Amii); a Canada CIFAR AI Chair, Amii; Digital Research Alliance of Canada; Mitacs; and the National Science and Engineering Research Council.
Ethics Statement
The goal of preference and scalar-based RL is to learn reward functions that encode human preferences. However, this necessitates interaction with human users, raising concerns about potential biases in the collected preferences. If these preferences are primarily drawn from a non-representative group, the resulting RL system may unfairly prioritize the desires or needs of that group over others. To that end, we performed a human subject study that was reviewed and approved by an external ethics committee. This ensured the responsible and ethical collection of human preferences. In addition, we took care to gather participants from a wide range of both demographic and educational backgrounds, which we detail in Section 5.3.
Reproducibility Statement
To ensure the reproducibility of our work, we provide a link to our code repository: https://github.com/cmuslima/SDP_ICLR. We also provide pseudocode in Section 4. Moreover, we outline extensive training and evaluation details in Section 5.1 and Appendix A.2. We also outline all hyperparameters used for all algorithms and environments in Appendix A.2, Tables 1-- 5. Lastly, for the human subject study, we provide details on the instructions provided to the participants in Appendix B as well as a description of the preference interface used. In our released codebase, we also provide the specific code files used to run the experiments in the user study. In addition, an anonymized version of the human subject data will be released, as stated in our ethics approval.
@@ Acknowledgments
Appendix A Simulated Experiment Details
@@ A.1 Benchmarks
MRN
MRN is a preference learning algorithm also integrated with PEBBLE. Unlike other PbRL algorithms, MRN is a bi-level optimization algorithm in which the actor and critic are updated in the inner loop, and the reward model is updated in the outer loop. Importantly, the reward model takes into account the performance of the critic on the preference data.
@@ A.2 Training Details
Furthermore, we use most of the existing reward model hyperparameters used in PEBBLE, however, we adjusted the following four hyperparameters: feedback frequency, amount of feedback per session, trajectory segment size (only for Meta-world), and activation function for the final NN layer. We adjusted the first two hyperparameters because PEBBLE originally used a significantly larger feedback budget, therefore we wanted the feedback schedule to better reflect a smaller feedback budget. We used a different trajectory segment size for Meta-world because we wanted to keep the segment sizes the same across both the DMControl and Meta-world environments. Moreover, we found that the output activation function could significantly affect learning, therefore we tested all benchmarks using both Tanh (original activation used) and Leaky-ReLU and chose the reward model that achieved better final performance. For the RUNE and SURF baselines, we use any hyperparameters associated with their specific algorithm according to the original paper (see Table 2). For a fair comparison with SDP, we provide all human-in-the-loop baselines (e.g., PEBBLE, R-PEBBLE, Deep TAMER, RUNE, and SURF) with the sub-optimal data set to be used in both the reward model and by the RL agent.
Furthermore, to select trajectory segments for the teacher to provide feedback on, we use uniform sampling in the DMControl tasks and disagreement sampling in the Meta-world tasks. Disagreement sampling is a popular active learning approach in which trajectories with higher uncertainty (based on an ensemble of neural networks) are more likely to be sampled (Christiano et al., 2017). As for the SAC hyperparameters, we use the values found in Tables 3-4.
@@
Segment Size 50
Random Steps
(i.e., Sub-optimal Data Transitions) 50000
Unsupervised Exploration Steps 9000
Consider Unsupervised Explorations
as Sub-optimal False (Walker-walk, Cartpole-swingup, Quadruped-walk – PbRL)
True (others)
Frequency of Feedback 20000 (DMControl)
10000 (Window-open, Door-unlock, Door-lock)
5000 (Hammer, Drawer-open)
Feedback Budgets in Ablations 2000 (Door-open)
1000 (Cheetah-run, Walker-walk)
Feedback Queries per Session 50 (Hammer, Drawer-open)
8 (Cartpole-swingup)
20 (others)
Sampling Scheme Disagreement Sampling (Metaworld)
Uniform Sampling (DMControl)
Training Epochs 200 (Window-open, SURF and SDP + SURF)
50 (others)
Learning Rate 3 ×$\times 10^{-4}$
Intermediate Neural Network Activation Leaky ReLU
Batch Size 128
Hidden Layers 4
Neurons per Hidden Layer 128
Loss Function Mean Squared Error (Scalar Feedback)
Cross Entropy Loss (Preference Feedback)
@@
For Reward Model Pre-training Phase in SDP
Last Layer Neural Network Activation Tanh
For Human-in-the-Loop Algorithm
Last Layer Neural Network Activation Leaky ReLU (PEBBLE, RUNE, SURF)
Tanh (SDP – Hammer, Drawer-open, Walker-walk)
Leaky ReLU (SDP – others)
Hyperparameter Value
Specific RUNE Hyperparameters
Beta Schedule Linear Decay
Beta Init 0.05
Beta Decay $10^{-5}$
Specific SURF Hyperparameters
Threshold $\lambda$ 0.99
Unlabeled Batch Ratio 4
Loss Weight 1
Min/Max Length of Cropped Segment [45, 55]
Segment Length Before Cropping 60
Specific MRN Hyperparameters
Meta Steps 1000 (Walker-walk)
3000 (Quadruped-walk)
10000 (Door-open)
5000 (others)
Table 2: Baseline Specific Hyperparameters.
Hyperparameter Value
Optimizer Adam (Kingma & Ba, 2015)
Discount 0.99
Alpha Learning Rate $10^{-4}$
Actor Betas 0.9, 0.999
Critic Betas 0.9, 0.999
Alpha Betas 0.9, 0.999
Target Smoothing Coefficient 0.005
Actor Update Frequency 1
@@
Init Temperature 0.1
Network Type MLP
Nonlinearity ReLU
Gradient Updates per Step 1
Table 3: Hyperparameters for SAC that were shared by all algorithms. SAC is the standard RL algorithm that learns from the environment’s true reward signal, not via preference learning.
@@
DMControl
Batch Size 512 (Cartpole-swingup)
1024 (others)
Hidden Layers 2
Neurons per Hidden Layer 256 (Cartpole-swingup)
1024 (others)
Actor/Critic Learning Rate $5\times 10^{-5}$ (Cheetah-run preference feedback)
$10^{-4}$ (Cheetah-run scalar feedback)
$5\times 10^{-4}$ (Walker-walk)
$10^{-4}$ (others)
Training Steps $0.5\times 10^{6}$ (Walker-walk, Cartpole-swingup)
$10^{6}$ (others)
Metaworld
Batch Size 512
Hidden Layers 3
Neurons per Hidden Layer 256
Actor/Critic Learning Rate $3\times 10^{-4}$
Training Steps $0.5\times 10^{6}$ (Door-lock and Window-open)
$10^{6}$ (others)
Table 4: Specific SAC hyperparameters that were tuned for the DMControl and Metaworld experiments. The majority of these hyperparameters were selected from the PEBBLE repo (Lee et al., 2021a).
@@
SAC
Training Steps $0.3\times 10^{6}$
Batch Size 1024
Hidden Layers 2
Neurons per Hidden Layer 1024
Actor/Critic Learning Rate $10^{-4}$
Reward Model
Consider Unsupervised Explorations as Sub-optimal False
Frequency of Feedback 20000
Feedback Queries per Session 10 (Pendulum-swingup)
8 (Cartpole-swingup)
Sampling Scheme Uniform Sampling